# On Scaling Up 3D Gaussian Splatting Training

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) is increasingly popular for 3D reconstruction due to its superior visual quality and rendering speed. However, 3DGS training currently occurs on a single GPU, limiting its ability to handle high-resolution and large-scale 3D reconstruction tasks due to memory constraints. We introduce Grendel, a distributed system designed to partition 3DGS parameters and parallelize computation across multiple GPUs. As each Gaussian affects a small, dynamic subset of rendered pixels, Grendel employs sparse all-to-all communication to transfer the necessary Gaussians to pixel partitions and performs dynamic load balancing. Unlike existing 3DGS systems that train using one camera view image at a time, Grendel supports batched training with multiple views. We explore various optimization hyperparameter scaling strategies and find that a simple sqrt(batch_size) scaling rule is highly effective. Evaluations using large-scale, high-resolution scenes show that Grendel enhances rendering quality by scaling up 3DGS parameters across multiple GPUs. On the “Rubble” dataset, we achieve a test PSNR of 27.28 by distributing 40.4 million Gaussians across 16 GPUs, compared to a PSNR of 26.28 using 11.2 million Gaussians on a single GPU. Grendel is an open-source project available at: https://github.com/nyu-systems/Grendel-GS
Abstract
3D Gaussian Splatting (3DGS) is gaining popularity for 3D reconstruction because of its superior visual quality and rendering speed. However, training is currently done on a single GPU, and thus cannot handle high-resolution and large-scale 3D reconstruction tasks due to the GPU’s memory capacity limit. We build a distributed system, called Grendel, to partition 3DGS’ parameters and parallelize its computation across multiple GPUs. As each Gaussian affects a small and changing subset of rendered pixels, Grendel relies on sparse all-to-all communication to transfer each required Gaussian to a pixel partition and performs dynamic load balancing. Unlike existing 3DGS systems that train using one camera view image at a time, Grendel supports batched training using multiple views. We explore different learning rate scaling strategies and identify the simple sqrt(batch_size) scaling rule to be highly effective. Evaluation using large-scale high-resolution scenes show that Grendel can improve rendering quality by scaling up 3DGS quantity using multiple GPUs. On the “Rubble” dataset, we achieve a test PSNR of 27.28 by distributing 40.4 million Gaussians across 16 GPUs. By comparison, one achieves a PSNR of 26.28 when using 11.2 million Gaussians in order to fit in a single GPU’s memory.
@@ 1 Introduction
3D Gaussian Splatting (3DGS)[8] has emerged as a popular technique for 3D novel view synthesis, primarily due to its faster training and rendering compared to previous approaches such as NeRF [21]. However, most existing 3DGS pipelines are constrained to using a single GPU for training, creating memory and computation bottlenecks when applied to high-resolution or larger-scale scenes. For example, the standard Rubble dataset [29] contains 1657 images, each with a 4K resolution. A single A100 40GB GPU can hold up to 11.2 million Gaussians – well below the quality saturation point for 3DGS. As we demonstrate in Section 5.2.1, increasing the number of Gaussians continues to improve reconstruction quality. Therefore, in order to scale up 3DGS training in terms of parameter count and speed, we develop the Grendel distributed training system, consisting of distributing 3DGS training across multiple GPUs and an empirical rule to automatically adapt training hyperparameters based on the batch size.
Current approaches that work around the single-device 3DGS training pipeline limitations when handling large-scale scenes typically involve compression schemes or selective rendering [9, 27, 19]. However, these methods are still confined to a single GPU and are fundamentally limited by similar memory and computational bottlenecks when scaling up further.
As large-scale distributed machine learning has become industry standard and a cornerstone for many state-of-the-art AI models such as Large Language Models [28, 26], it is reasonable to expect that distributing workloads across multiple GPUs leads to similar performance and scale improvements, where it allows 3DGS to handle larger, higher-resolution scenes and produce higher-quality outputs.
However, 3DGS stands out as a gradient-based but not a neural network-based training framework. It features a unique computation pipeline with dynamic and imbalanced workload patterns within a training iteration. Consequently, existing SysML frameworks and toolkits, such as those in [28, 26, 14, 34], do not apply to this domain. The frameworks that target neural networks assume a consistent and balanced workload consisting of a series of regular tensor operations (e.g. GEMM) that are distinct from the dynamic, non-uniform computation patterns exhibited by 3DGS training and rendering. Therefore, designing a scalable distributed system for 3DGS is uniquely challenging.
@@
In this paper, we present several key observations on scaling up Gaussian Splatting training that inform the design of our distributed learning pipeline for 3DGS. For instance, we note that each stage of the 3DGS training pipeline in an iteration can be effectively parallelized, but the axes of parallelization differ across stages. In the mixed parallelism of the 3DGS training pipeline, some computations operate on individual output pixels (allowing for pixel-wise parallelism), while others operate on individual 3D Gaussians (allowing for Gaussian-wise parallelism). This enables parallelism but necessitates communication between pipeline stages. To minimize communication, we also observe that 3DGS exhibits spatial locality, where only a small number of Gaussians affect each image patch. Finally, the computational intensity of rendering an output pixel changes as training progresses, so we identify the dynamic and unbalanced workloads. As a result, any static workload partitioning strategy will be suboptimal.
@@
In this paper, we describe Grendel, a distributed 3DGS training framework designed under these observations. Grendel uses Gaussian-wise distribution–that is, it distributes Gaussians across GPUs–for steps in a training iteration that exhibit Gaussian-wise parallelism, and pixel-wise distribution for other steps. It minimizes the communication overhead when switching between Gaussian-wise and pixel-wise distribution by assigning contiguous image areas to GPUs during pixel-wise distribution and exploiting spatial locality to minimize the number of Gaussians transferred among GPUs. Finally, Grendel employs a dynamic load balancer that uses previous training iterations to distribute pixel-wise computations to minimize workload imbalance.
@@
Grendel additionally scales up training by batching up multiple images. This differs from conventional 3DGS training that exclusively uses a batch size of 1, which would lead to reduced GPU utilization in our distributed framework. To maintain data efficiency and reconstruction quality with larger batches, one needs to re-tune optimizer hyperparameters. To this end, we introduce an automatic hyperparameter scaling rule for batched 3DGS training based on a heuristical independent gradients hypothesis. We empirically validate the effectiveness of our proposed approach — Grendel supports distributed training with large batch sizes (we test up to 32) while maintaining reconstruction quality and data efficiency compared to batch size $=1$.
@@
  • We show that Grendel enables high-resolution large scale scene rendering: we use 16 GPUs and render 4K images for large-scale Rubble scene from MegaNERF [29]. For this scene, Grendel uses 40.4 million Gaussians to achieve a PSNR of 27.28, outperforming the current state-of-the-art. The memory required exceeds a single GPU’s capacity, making it difficult to render this scene at this quality without Grendel’s techniques.
@@ 2 Gaussian Splatting: Background, Opportunities and Challenges
3D Gaussian Splatting [8] (3DGS) is a rendering method that represents 3D scenes using a (potentially large) set of anistropic 3D Gaussians. Each 3D Gaussian is represented by four learnable parameters: (a) its 3D position $x_{i}\in\mathbb{R}^{3}$; (b) its shape described by a 3D covariance matrix computed using the Guassian’s scaling vector $s_{i}\in\mathbb{R}^{3}$ and rotation vector $q_{i}\in\mathbb{R}^{4}$; (c) its opacity $\alpha_{i}\in\mathbb{R}$; and (d) its spherical harmonics $sh_{i}\in\mathbb{R}^{48}$. The color contribution of each Gaussian is detemined by these parameters and by the viewing-direction.
@@ 2.1 Background on 3D Gaussian Training
To train 3DGS, the user provides an initial point cloud (may be random or estimated) for a scene and a set of posed images from different angles. The training process initializes Gaussians using the point cloud. Each training step selects a random camera view and uses the current Gaussian parameters to render the view. It then computes loss by comparing the rendered image to the ground truth, and uses back-propagation to update the Gaussian parameters. The training process also uses an adaptive densification mechanism to add Gaussians to under-reconstructed areas. Gaussians are added by cloning or splitting existing ones based on their position variance and scale threshold, with more details in A.1.
@@ 2.2 Opportunities and Challenges in distributing 3DGS
Opportunity: mixed parallelism.
Each of the steps described above is inherently parallel but requires different kinds of work partitioning. In particular, the Gaussian transformation step operates on individual Gaussians and thus should be partitioned by Gaussians. On the other hand, the rendering and loss calculation steps operate on individual pixels(or pixels windows for SSIM loss) and thus should be partitioned by pixel.
Opportunity: spatial locality.
Most Gaussians intersect a small contiguous area of the rendered image due to their typically small radius. As illustrated in Figure 3, 90% of the 3D Gaussians in three scenes (Rubble, Bicycle, and Train) have a radius $<2\%\leavevmode\nobreak\ \text{of image width}$. Consequently, a pixel is affected by a small subset of the scene’s 3D Gaussians, with significant overlap among neighboring pixels’ Gaussians.
Figure 3: The footprint radius of most Guassians is small relative to the image size.
Challenge: dynamic and unbalanced workloads.
Different image areas intersect varying quantities of Gaussians. For instance, an image region containing the sky likely corresponds to fewer Gaussians than a region with a person. Additionally, the density, position, shape, and opacity of Gaussians change throughout training. Therefore, the number of Gaussians and their mapping to pixels evolve over time, leading to computational workload imbalances across different image regions and over the training period. Fixed partitioning schemes thus suffer from load imbalance.
Challenge: absence of batching.
Current 3DGS systems process images one at a time, which suffices for single GPU training. However, as shown in §5, this approach is inefficient in a distributed setting with multiple GPUs. Effective training with larger batch sizes necessitates an understanding of the unique optimization dynamics of 3DGS, which may differ from those of conventional neural networks.
@@ 3.1 Mixed parallelism training
Pixel-wise Distribution. We distribute contiguous image areas across GPUs for the image rendering and loss computation steps. Distributing contiguous areas allows us to exploit spatial locality and reduce the number of Gaussians transferred among GPUs in this step. In our implementation, we partition each image in a batch by dividing it into $16\times 16$-pixel blocks, serializing the blocks, and then distributing consecutive subsequences of blocks to different GPUs using an adaptive strategy (§3.2). For batching, each GPU can be assigned blocks from different images in a batch, as shown in Figure 2(b).
Figure 4: Each GPU only considers Gaussians whose footprints intersect with its assigned pixel render area. Every square represents a 16 × 16 block of pixels.
@@
Although Grendel’s design bears some resemblance to FSDP [34] used for distributed deep neural network training, there are some important differences. Firstly, unlike weight sharding in FSDP, Gaussian-wise distribution in Grendel is not merely for storage but for also for computation (the Gaussian transformation). Secondly, unlike FSDP which transfers weight shards using the dense all-gather communication, Grendel transfers only relevant Guassians using sparse all-to-all communication.
@@ 3.2 Iterative Workload Rebalancing
Pixel-wise Distribution Rebalancing.
As discussed in §2.2, the computational load of rendering a pixel varies across space (different pixels) and time (different training iterations). Thus, unlike in distributed neural network training, a uniform or fixed distribution cannot guarantee balanced workloads, so an adaptive pixel distribution strategy is needed.
@@
1:$ET_{j}$, $B$ (number of pixel blocks), $G$ (number of GPUs)
2:$DP$ (division points)
3:$CT\leftarrow\textsc{torch.cumsum}(ET)$triangleright Cumulative sum of ET
4:$ET_{gpu}\leftarrow CT[B-1]/G$triangleright Estimated runtime per GPU
5:$TH\leftarrow\textsc{torch.arange}(0,G)\cdot ET_{gpu}$triangleright Thresholds for Division Points
6:$DP\leftarrow\textsc{torch.searchsorted}(CT,TH)$triangleright Division Points
7:return $DP$
Gaussian-wise Distribution Rebalancing.
When training starts, we distribute 3D Gaussians uniformly among the GPUs. As training progresses, new Gaussians are added by cloning and splitting existing ones(§2.1). Newly added Gaussians make the distribution imbalanced as different Gaussians densify at different rates that depend on the scene’s local details. Therefore, we redistribute the 3D Gaussians after every few densification steps to restore uniformity.
@@ 4 Scaling Hyperparameters for Batched Training
To efficiently scale to multiple GPUs, Grendel increases the batch size beyond one so it can partition training into a batch of images in addition to pixels inside each image, as shown in Figure 2(b).
However, increasing the batch size without adjusting hyperparameters, particularly the learning rate, can result in unstable and inefficient training [4, 25], and hyperparameter tuning is a time-consuming and tedious process. Though there have been works that simplify learning-rate tuning for training deep neural networks, their settings are distinct from ours since they either build on SGD [4] (where we use Adam) or they leverage the layer-wise structure of neural networks [3, 32]. We propose an automatic hyperparameter scaling rule for the Adam optimizer to enable hyperparameter-tuning-free batched 3DGS training. Our result is driven by the Independent Gradients Hypothesis unique to 3DGS. Interestingly, our proposed square-root scaling rule coincides with recent works that study Adam learning rate scaling for neural network training even though they make distinct assumptions [5, 13].
@@
Independent Gradients Hypothesis. To arrive at these scaling rules, it is helpful to consider 3D GS training in a simplified setting. Suppose that the gradients received by 3D Gaussian parameters from each camera view are independent of gradients induced by other views. Consequently, if we are given a batch of $b$ camera views, taking $b$ sequential gradient descent steps for each view in the batch is equivalent to taking one bigger step where the gradients are summed together. If we were using the vanilla gradient descent algorithm and averaging the gradients in a batch, setting the learning rate to scale linearly with the batch size achieves this equivalence. However, 3D GS uses Adam, an adaptive learning rate optimizer that (1) divides the gradients by the square root of the per-parameter second moment estimate, and (2) uses momentum to combine current gradients and past gradients in an exponential-moving-average fashion, making a bigger update different from simply summing up smaller batch-size-one updates. Under the independent gradients hypothesis, we derive the following corrections to Adam hyperparameters to approximate batch-size-one training with a larger batch:
@@
Thus, setting the learning rate $\lambda^{\prime}=\lambda\times\sqrt{b}$ allows the batch update $\Delta^{\{B\}}$ to match with the total individual updates $\sum_{k\in B}\Delta^{\{k\}}$.
Alongside the square-root learning rate scaling (Eq 1), we also propose an exponential momentum scaling to accommodate larger batches (Eq 2). Initially used by [2], this rule scales the momentum parameters with $\beta^{\prime}=\beta^{\text{batch\_size}}$, which exponentially decreases the influence of past gradients when the batch size increases. We wish to stress that in the real world, even though some cameras share similar poses, a set of random cameras generally observe different parts of a scene, hence the gradients in a batch are mostly sparse and can be thought of as roughly independent. We empirically study the independent gradient hypothesis and evaluate our proposed scaling rules.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) is increasingly popular for 3D reconstruction due to its superior visual quality and rendering speed. However, 3DGS training currently occurs on a single GPU, limiting its ability to handle high-resolution and large-scale 3D reconstruction tasks due to memory constraints. We introduce Grendel, a distributed system designed to partition 3DGS parameters and parallelize computation across multiple GPUs. As each Gaussian affects a small, dynamic subset of rendered pixels, Grendel employs sparse all-to-all communication to transfer the necessary Gaussians to pixel partitions and performs dynamic load balancing. Unlike existing 3DGS systems that train using one camera view image at a time, Grendel supports batched training with multiple views. We explore various optimization hyperparameter scaling strategies and find that a simple sqrt(batch_size) scaling rule is highly effective. Evaluations using large-scale, high-resolution scenes show that Grendel enhances rendering quality by scaling up 3DGS parameters across multiple GPUs. On the 4K “Rubble” dataset, we achieve a test PSNR of 27.28 by distributing 40.4 million Gaussians across 16 GPUs, compared to a PSNR of 26.28 using 11.2 million Gaussians on a single GPU. Grendel is an open-source project available at: https://github.com/nyu-systems/Grendel-GS
@@ 1 Introduction
3D Gaussian Splatting (Kerbl et al., 2023) (3DGS) has emerged as a popular technique for 3D novel view synthesis, primarily due to its faster training and rendering compared to previous approaches such as NeRF (Mildenhall et al., 2020). However, most existing 3DGS pipelines are constrained to using a single GPU for training, creating memory and computation bottlenecks when applied to high-resolution or larger-scale scenes. For example, the standard Rubble dataset (Turki et al., 2022) contains 1657 images, each with a 4K resolution. A single A100 40GB GPU can hold up to 11.2 million Gaussians – well below the quality saturation point for 3DGS. As we demonstrate in Section 5.2.1, increasing the number of Gaussians continues to improve reconstruction quality. Therefore, in order to scale up 3DGS training in terms of parameter count and speed, we develop the Grendel system, which distributes 3DGS training across multiple GPUs and uses an empirical rule to automatically adapt training hyperparameters based on the batch size.
Existing work on scaling 3DGS to large-scale scenes have explored approaches such as divide-and-conquer Lin et al. (2024); Liu et al. (2024); Kerbl et al. (2024), level-of-detail representation and rendering (Kerbl et al., 2024; Ren et al., 2024; Lu et al., 2024). These methods are complementary to the system-level parallelization approach that we adopt to scale 3DGS beyond the memory and compute limitations of a single GPU.
As distributed training frameworks have become widely used for many state-of-the-art DNN models such as LLMs (Shoeybi et al., 2020; Rajbhandari et al., 2020), it is tempting to also use them to distribute 3DGS. However, although 3DGS uses gradient-based optimization, it is not based on neural networks. Specifically, it features a unique computation pipeline with dynamic and imbalanced workload patterns. Consequently, existing DNN training frameworks (Shoeybi et al., 2020; Rajbhandari et al., 2020; Li et al., 2020; Zhao et al., 2023), which assume consistent and balanced workload with regular dense tensor operations, would not work well for 3DGS.
@@
In this paper, we present several key observations on scaling up 3DGS that inform the design of our distributed training pipeline. For instance, we note that each stage of the 3DGS training pipeline in an iteration can be effectively parallelized, but the axes of parallelization differ across stages, resulting in mixed parallelism. More concretely, in 3DGS, some computations operate on individual output pixels (allowing for pixel-wise parallelism), while others operate on individual 3D Gaussians (allowing for Gaussian-wise parallelism). Mixed parallelism necessitates data shuffling between stages. To minimize communication, we also observe that 3DGS exhibits spatial locality, where only a small number of Gaussians affect the rendering of each output image patch. Finally, the computational intensity of rendering an output pixel changes as training progresses. Such dynamic and unbalanced workloads will cause any static workload partitioning strategy to become suboptimal.
@@
In this paper, we describe Grendel, a distributed 3DGS training framework designed to leverage our above observations. Grendel uses Gaussian-wise distribution–that is, it distributes Gaussians across GPUs–for steps in a training iteration that exhibit Gaussian-wise parallelism, and pixel-wise distribution for other steps. It minimizes the communication overhead when switching between Gaussian-wise and pixel-wise distribution by assigning contiguous image areas to GPUs during pixel-wise distribution and exploiting spatial locality to minimize the number of Gaussians transferred among GPUs. Finally, Grendel employs a dynamic load balancer that uses previous training iterations to distribute pixel-wise computations to minimize workload imbalance.
@@
Grendel additionally scales up training by batching multiple images. This differs from conventional 3DGS training that exclusively uses a batch size of 1, which would lead to reduced GPU utilization in our distributed framework. To maintain data efficiency and reconstruction quality with larger batches, one needs to re-tune optimizer hyperparameters. To this end, we introduce an automatic hyperparameter scaling rule for batched 3DGS training based on a heuristical independent gradients hypothesis. We empirically validate the effectiveness of our proposed approach — Grendel supports distributed training with large batch sizes (we test up to 32) while maintaining reconstruction quality and data efficiency compared to batch size =1=1.
@@
  • We show that Grendel enables high-resolution large scale scene rendering: we use 16 GPUs and render 4K images for large-scale Rubble scene from MegaNERF (Turki et al., 2022). For this scene, Grendel uses 40.4 million Gaussians to achieve a PSNR of 27.28, outperforming the current state-of-the-art. The memory required exceeds a single GPU’s capacity, making it difficult to render this scene at this quality without Grendel’s techniques.
@@ 2 Gaussian Splatting: Background, Opportunities and Challenges
3D Gaussian Splatting (Kerbl et al., 2023) (3DGS) is a rendering method that represents 3D scenes using a (potentially large) set of anistropic 3D Gaussians. Each 3D Gaussian is represented by four learnable parameters: (a) its 3D position $x_{i}\in\mathbb{R}^{3}$; (b) its shape described by a 3D covariance matrix computed using the Guassian’s scaling vector $s_{i}\in\mathbb{R}^{3}$ and rotation vector $q_{i}\in\mathbb{R}^{4}$; (c) its opacity $\alpha_{i}\in\mathbb{R}$; and (d) its spherical harmonics $sh_{i}\in\mathbb{R}^{48}$. The color contribution of each Gaussian is detemined by these parameters and by the viewing-direction.
@@ 2.1 Background on 3D Gaussian Training
To train 3DGS, the user provides an initial point cloud (may be random or estimated) for a scene and a set of posed images from different angles. The training process initializes Gaussians using the point cloud. Each training step selects a random camera view and uses the current Gaussian parameters to render the view. It then computes loss by comparing the rendered image to the ground truth, and uses back-propagation to update the Gaussian parameters. The training process also uses an adaptive densification mechanism to add Gaussians to under-reconstructed areas, by cloning or splitting existing ones based on their position variance and scale threshold, with more details in A.1.
@@ 2.2 Opportunities and Challenges in distributing 3DGS
Opportunity: mixed parallelism. Each of the steps described above is inherently parallel but requires different kinds of work partitioning. In particular, the Gaussian transformation step operates on individual Gaussians and thus should be partitioned by Gaussians. On the other hand, the rendering and loss calculation steps operate on individual pixels(or pixels windows for SSIM loss) and thus should be partitioned by pixel.
Opportunity: spatial locality. Most Gaussians intersect a small contiguous area of the rendered image due to their typically small radius. As illustrated in Figure 3, 90% of the 3D Gaussians in three scenes (Rubble, Bicycle, and Train) have a radius $2\%~\text{of image width}$. Consequently, a pixel is affected by a small subset of the scene’s 3D Gaussians, with significant overlap among neighboring pixels’ Gaussians.
Figure 3: Cumulative percents of $Radius_{i}$ relative to image width
Challenge: dynamic and unbalanced workloads. Different image areas intersect varying quantities of Gaussians, as shown in Fig 5. For instance, an image region containing the sky likely corresponds to fewer Gaussians than a region with a person. Additionally, the density, position, shape, and opacity of Gaussians change throughout training. Therefore, the number of Gaussians and their mapping to pixels evolve over time, leading to computational workload imbalances across different image regions and over the training period. Fixed partitioning schemes thus suffer from load imbalance. Refer to the Appendix §A.4 for further details.
Challenge: absence of batching. Current 3DGS systems process images one at a time, which suffices for single GPU training. However, as shown in §5, this approach is inefficient in a distributed setting with multiple GPUs. Effective training with larger batch sizes necessitates an understanding of the unique optimization dynamics of 3DGS, which may differ from those of conventional neural networks.
@@ 3.1 Mixed parallelism training
Pixel-wise Distribution. We distribute contiguous image areas across GPUs for the image rendering and loss computation steps. Distributing contiguous areas allows us to exploit spatial locality and reduce the number of Gaussians transferred among GPUs. In our implementation, we partition each image in a batch by dividing it into $16\times 16$-pixel blocks, serializing the blocks, and then distributing consecutive subsequences of blocks to different GPUs using an adaptive strategy (§3.2). For batching, each GPU can be assigned blocks from different images in a batch, as shown in Figure 2(b).
Figure 4: We present a heatmap of the per-tile imbalance in the number of rendered Gaussians on the Train Dataset (Knapitsch et al., 2017). Redder tiles indicate more Gaussian splats. Distant, low-detail areas like the sky need fewer Gaussians than detailed foreground regions like the train, highlighting an imbalance in rendering intensity.
Figure 5: Each GPU only considers Gaussians whose footprints intersect with its assigned pixel render area.
@@
Although Grendel’s design bears some resemblance to FSDP (Zhao et al., 2023) used for distributed neural network training, there are important differences. Firstly, unlike weight sharding in FSDP, Gaussian-wise distribution in Grendel is not merely for storage but for also for computation (the Gaussian transformation). Secondly, unlike FSDP which transfers weight shards using the dense all-gather communication, Grendel transfers only relevant Guassians using sparse all-to-all communication.
@@ 3.2 Iterative Workload Rebalancing
Pixel-wise Distribution Rebalancing. As discussed in §2.2, the computational load of rendering a pixel varies across space (different pixels) and time (different training iterations). Thus, unlike in distributed neural network training, a uniform or fixed distribution cannot guarantee balanced workloads, so an adaptive pixel distribution strategy is needed.
@@
1:BB(number of pixel blocks), GG(number of GPUs), $ET_{j}$(Estimated runtime per pixel block)
2:DPDP (division points)
[…]
7:return DPDP
Gaussian-wise Distribution Rebalancing. When training starts, we distribute 3D Gaussians uniformly among the GPUs. As training progresses, new Gaussians are added by cloning and splitting existing ones(§2.1). Newly added Gaussians make the distribution imbalanced as different Gaussians densify at different rates that depend on the scene’s local details. Therefore, we redistribute the 3D Gaussians after every few densification steps to restore uniformity.
@@ 4 Scaling Hyperparameters for Batched Training
To efficiently scale to multiple GPUs, Grendel increases the batch size beyond one, enabling partitioning of both images and pixels within each image, as shown in Figure 2(b).
However, increasing the batch size without adjusting hyperparameters, particularly the learning rate, can result in unstable and inefficient training (Goyal et al., 2017; Qiao et al., 2021), and hyperparameter tuning is often tedious. Though some methods simplify learning-rate tuning for deep neural networks, they either build on SGD (Goyal et al., 2017) (we use Adam) or they leverage the layer-wise structure of neural networks (Ginsburg et al., 2018; You et al., 2020) (3DGS is not neural network). Our result is driven by the Independent Gradients Hypothesis for 3DGS training. Inspired by (Malladi et al., 2022), we derive a scaling rule for the hyperparameters of Adam, which suggests the same learning rate scaling as recent works (Malladi et al., 2022; Granziol et al., 2022) but a different scaling for $\beta_{1}$ and $\beta_{2}$ that works better for 3DGS.
@@
Independent Gradients Hypothesis. To derive these scaling rules, we first consider 3D GS training in a simplified setting, assuming that gradients calculated from each camera view are independent of those induced from other views. Consequently, if we are given a batch of $b$ camera views, taking $b$ sequential gradient descent steps for each view in the batch is equivalent to taking one bigger step where the gradients are summed together. If we were using the vanilla gradient descent algorithm and averaging the gradients in a batch, setting the learning rate to scale linearly with the batch size achieves this equivalence. However, 3D GS uses Adam, an adaptive learning rate optimizer that (1) divides the gradients by the square root of the per-parameter second moment estimate, and (2) uses momentum to combine current gradients and past gradients in an exponential-moving-average fashion, making a bigger update different from simply summing up smaller batch-size-one updates. Under the independent gradients hypothesis, we derive the following corrections to Adam hyperparameters to approximate batch-size-one training with a larger batch:
@@
Thus, setting the learning rate $\lambda^{\prime}=\lambda\times\sqrt{b}$ allows the batch update $\Delta^{\{B\}}$ to match with the total individual updates $\sum_{k\in B}\Delta^{\{k\}}$. Alongside the square-root learning rate scaling (Eq 1), we also propose an exponential momentum scaling to accommodate larger batches (Eq 2). Initially used by Busbridge et al. (2023), this rule scales the momentum parameters with $\beta^{\prime}=\beta^{\text{batch\_size}}$, which exponentially decreases the influence of past gradients when the batch size increases.
We wish to stress that in the real world, even though some cameras share similar poses, a set of random cameras generally observe different parts of a scene, hence the gradients in a batch are mostly sparse and can be thought of as roughly independent. We empirically study the independent gradient hypothesis and evaluate our proposed scaling rules.